# Splat Feature Solver

### v1

@@ Abstract
Feature lifting has emerged as a crucial component in 3D scene understanding, enabling the attachment of rich image feature descriptors (e.g., DINO, CLIP) onto splat-based 3D representations. The core challenge lies in optimally assigning rich general attributes to 3D primitives while addressing the inconsistency issues from multi-view images. We present a unified, kernel- and feature-agnostic formulation of the feature lifting problem as a sparse linear inverse problem, which can be solved efficiently in closed form. Our approach admits a provable upper bound on the global optimal error under convex losses for delivering high quality lifted features. To address inconsistencies and noise in multi-view observations, we introduce two complementary regularization strategies to stabilize the solution and enhance semantic fidelity. Tikhonov Guidance enforces numerical stability through soft diagonal dominance, while Post-Lifting Aggregation filters noisy inputs via feature clustering. Extensive experiments demonstrate that our approach achieves state-of-the-art performance on open-vocabulary 3D segmentation benchmarks, outperforming training-based, grouping-based, and heuristic-forward baselines while producing the lifted features in minutes. Code is available at github. We also have a webpage
Figure 1: Overview of Our Framework. Our method comprises three main components: Input, Solver, and Output. The Inputs (left column) include: (1) a precomputed general splat representation; (2) sensor parameters (camera intrinsics, extrinsics, and image resolution); and (3) dense feature observations, such as MaskCLIP or DINO embeddings. These inputs feed into our Feature Lifting Solver (central gray box), which formulates the feature lifting equation. Here, the splat representation and sensor parameters produce the Splats Sensor Matrix, while observations form the equation’s right-hand side (both highlighted in red). To solve for the feature parameters (also highlighted in red), we enhance numerical stability by polarizing alpha values to encourage diagonal dominance, and employ Post-Lifting Aggregation to filter out incomplete or noisy masks. The Output (upper-right blue box) consists of the lifted feature parameters, supporting upstream tasks (e.g., clustering, PCA visualization) and downstream applications (lower-right blue box) like 3D localization and semantic segmentation.
Introduction
@@
Recent advances in splat representations—such as 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023), 2D Gaussian Splatting (Huang et al. 2024), and Deformable Beta Splatting (Liu et al. 2025)—have enabled real-time, high-fidelity scene rendering by modeling geometry with compact, explicit primitives. These splat-based methods combine differentiable projection, alpha compositing, and efficient visibility-aware rasterization to preserve geometric detail and multi-view consistency, supporting applications ranging from style transfer (Liu et al. 2024; Galerne et al. 2025) to scene understanding (Guo et al. 2024; Qin et al. 2024; Shi et al. 2024; Jun-Seong et al. 2025; Cheng et al. 2024; Cen et al. 2025; Dou et al. 2024; Gu et al. 2024; Zuo et al. 2025; Peng et al. 2024b). Nevertheless, enriching these primitives with detailed descriptors—such as CLIP, DINO, General ViT, and CNN features—remains challenging due to the inherent difficulty of coherently lifting 2D observations into consistent 3D representations.
@@
To address this challenge, recent efforts have focused on lifting per-pixel semantic descriptors (e.g., CLIP, DINO) onto 3D primitives, enabling semantic tasks such as segmentation and querying. Existing semantic feature lifting methods broadly fall into three categories: training-based optimization (Shi et al. 2024; Qin et al. 2024; Zhou et al. 2024; Qiu et al. 2024; Zuo et al. 2025), which pioneered in embedding semantic onto 3D primitives via multi-view training; grouping-based association (Wu et al. 2025; Peng et al. 2024b; Gu et al. 2024; Liang et al. 2024), which improved efficiency through feature clustering; and heuristic forward methods (Guo et al. 2024; Joseph, Amrutur, and Bhatnagar 2024; Dou et al. 2024; Chacko et al. 2025; Jun-Seong et al. 2025; Cheng et al. 2024), which prioritize speed and directly project semantic features onto 3D primitives.
@@
While existing methods have achieved promising results, a unified theoretical framework for feature lifting remains underexplored, especially concerning the heuristic forward methods that bypass the training process. Building upon prior efforts, we identify several open challenges that limit generalization and theoretical understanding. First, establishing a rigorous mathematical formulation for defining feature lifting could be beneficial to enable formal analysis, generalization, and optimization. Second, in the absence of a formal definition, existing approaches currently lack theoretical guarantees regarding the quality of the lifted features, leaving some uncertainty about their proximity to a globally optimal solution. Third, many existing methods are specialized for particular feature types or geometric kernels, which may limit generalization across broader settings. Finally, most prior works do not explicitly account for the inherently noisy nature of collected data, where multi-view inconsistencies and observational noise can introduce ambiguity. Our formulation accounts for these challenges through a theoretically grounded linear inverse framework.
Our key contributions are summarized as follows:
@@
  • Introduce the formulation of feature lifting as a sparse linear inverse problem, where each per-primitive descriptor is recovered via a globally consistent solution to the system $AX=B$
@@
The outline of our frame work is shown in Fig.1. Our formulation is the first to formalize the feature lifting problem as a general linear inverse problem, and can therefore be applied across diverse splat primitives (e.g., 3DGS, 2DGS, Beta Splats) and dense feature modalities (e.g., CLIP, DINO, ViT, ResNet features). Moreover, it establishes a new state-of-the-art on downstream tasks such as open-vocabulary 3D semantic segmentation, outperforming existing training-based, grouping-based, and heuristic-forward baselines in mIoUs.
Related Work
@@
Joint Training-based Methods—such as LangSplat (Qin et al. 2024), LeGaussian (Shi et al. 2024), FMGS (Zuo et al. 2025), and FeatureSplats (Qiu et al. 2024)—jointly optimize scene geometry and high‑dimensional feature embeddings to directly learn splat‑wise semantic descriptors. While effective, end-to-end optimization over large descriptor sets and primitives can be computationally intensive, often requiring significant memory and training time. To alleviate resource demands, these approaches employ descriptor compression (e.g., PCA in LangSplat (Qin et al. 2024), quantization in LeGaussian (Shi et al. 2024), and deep auto‑encoders in FeatureSplats (Qiu et al. 2024)) or reduce the number of primitives (FMGS (Zuo et al. 2025)). These strategies help manage resource usage, though they may introduce trade-offs in terms of geometric detail or feature precision.
@@
Grouping-based Methods first extract 2D region or instance masks using models such as SAM (Kirillov et al. 2023) and SAM2 (Ravi et al. 2024). Following that, each mask is linked to 3D Gaussians via a lightweight training process. Finally, features are aggregated into 3D either by additional optimization or by direct unprojection (Wu et al. 2025; Peng et al. 2024b; Gu et al. 2024; Liang et al. 2024; Cen et al. 2025). While these approaches reduce the cost of joint optimization, they still require one to two hours of scene‑specific optimization. Furthermore, these methods depend on SAM’s per-view masks, while powerful, are designed primarily for instance segmentation and may not be directly suited for noisy, dense, pixel-wise feature lifting.
@@
In particular, LAGA (Cen et al. 2025) uses per-view SAM masks (Kirillov et al. 2023) and trains an affinity model to cluster 3D Gaussian splats, subsequently associating splat clusters across views. However, this approach assumes that feature variance reflects true semantic differences and is therefore view-dependent, introducing a dynamic K-means clustering step that adds complexity and may increase memory and computational requirements. In contrast, our analysis suggests that such discrepancies often stem from mask inaccuracies—such as one mask isolating only the target while another over-segments and includes adjacent objects—rather than from genuine viewpoint changes. Our convex-regularized inverse problem solver suppresses this noise directly—without requiring additional training, multi-level clustering algorithms, or view-dependent features assigned to one primitive.
@@
Heuristic Forward Methods-e.g., Argmax Lifting (Chacko et al. 2025), Occam’s LGS (Cheng et al. 2024), Semantic Gaussian (Guo et al. 2024), gradient-guided splitting (Joseph, Amrutur, and Bhatnagar 2024), CosegGaussians (Dou et al. 2024), and DrSplats (Jun-Seong et al. 2025)—are analytic and training-free, offer high efficiency and simplicity. However, they often lack a formal mathematical foundation and sensitive to noisy inputs. Interestingly, three recent heuristic pipelines—CosegGaussians (Dou et al. 2024), Occam’s LGS (Cheng et al. 2024), and DrSplats (Jun-Seong et al. 2025)—independently proposed the identical row-sum weighting rule (row-sum preconditioner), underscoring its broad practical effectiveness. However, while CosegGaussians applied the row‐sum weighting scheme $x_{i}=\sum_{j}w_{i,j}b_{j}/\sum_{j}w_{i,j}$ , it did so without an accompanying theoretical interpretation. Occam’s LGS later offered a maximum‐likelihood interpretation of the same weighted‐average rule but provided no error bound. Dr. Splat simplifies the preconditioner further by summing only the top-$k$ contributions per primitive, trading off some theoretical rigor for computational efficiency.
@@
Feature Splat Solver
@@
Problem Definition
Splats Rendering
To justify formulating the Feature Lifting problem as a linear inverse problem, we first review the core splats rendering process. Most splats primitives, including planar 2D Gaussian Splats (Huang et al. 2024), volumetric 3D Gaussian Splats (Yu et al. 2024; Kerbl et al. 2023), and Deformable Beta Splats (Liu et al. 2025), employ a fixed, depth-sorted alpha-blending pipeline. Concretely, for each viewing ray $r$, there is a sequence of primitives ordered from front to back, with their respective contribution to the final color denoted as $\omega_{p}$. The background color is $C_{b}$. The rendered ray color $C_{r}$ is given below:
@@
The rendered color $C_{r}$ is usually very close (i.e. PSNR is more than 24 according to (Yu et al. 2024; Kerbl et al. 2023; Liu et al. 2025)) to the given observation $\hat{C}_{r}$ in the training data set.
The first property arises from the tile-based rendering design of Gaussian Splats(Kerbl et al. 2023). Since the tile size is typically set to $16\times 16$, most splats are excluded from rendering on any given tile.
The second property is the row stochastic property. It is less than one because the standard alpha‐blending heuristic of terminating ray integration once the composite opacity approaches one. It is close to one based on the assumption that scene geometry should block any further contribution from background color during training. In practice, we randomize the background color during training, forcing splats to fully occlude the randomly colored background from all viewing directions. Further justification of this property can be found in the appendix.F.
The third observation is noted in a recent splats training method (Yu et al. 2024; Huang et al. 2024; Kerbl et al. 2023; Liu et al. 2025). The aforementioned properties will be used to justify our linear inverse formulation.
Feature Lifting Equation
@@
Inverse Problem
First, we ask whether Equ.3 admits a solution, and, whether the given solution is unique. In typical settings—where $A\in R^{R\times P}$ with $R\gg P$— the system is overdetermined and generally inconsistent, meaning no exact $X$ satisfies (3). Instead, we consider the least squares formulation (or more generally, a convex loss minimization problem):
@@
where the minimizer exists for any $B$ and lies in the column space of $A$. In our context, Property 3 guarantees that real‐world observations $B$ , such as RGB color, approximately lie within $\mathrm{range}(A)$. As a result, the least-squares residual is small, and a near-exact lift exists. Intuitively, any observation derived from the same RGB signal should have a near-exact lift. From a semantic viewpoint, we also require a one‐to‐one mapping. Each geometric primitive should admit a single descriptor, mirroring an ideal embedding (e.g. a perfect 3D CLIP) that assigns one feature per object. Thus, uniqueness in (5) is not just for mathematical convenience, but for the purpose of aligning with downstream task requirements. Any deviation from uniqueness signals is either due to noisy inputs or suboptimal solvers—precisely the singular, and misleading-mask problem we address through our convex regularization and preconditioning strategies.
@@
Solver
In this session, we will introduce our solver under the well-posedness condition and give an optimal upper bound of our solver. In our setting, the least‐squares problem in Eq. (5) is convex, so stochastic methods such as SGD provably converge to the global minimizer (Esser, Zhang, and Chan 2010; Saad 2003). However, cold‐start training can be prohibitively slow. Prior works (e.g. (Chacko et al. 2025; Dou et al. 2024; Jun-Seong et al. 2025; Cheng et al. 2024; Joseph, Amrutur, and Bhatnagar 2024)) have explored efficient one-shot approximations that show practical value but do not offer formal guarantees. To address this, we first introduce the row‐sum preconditioner as a principled, closed‐form solver. We then develop an analytical framework that (i) quantifies its $(1+\beta)$-approximation error under both L2 and general convex losses, and (ii) interprets the behavior of existing heuristics as special cases.
$D^{\frac{1}{2}}(A^{T}A)$ $=\sqrt{D(A^{T}A)}$ (6)
$x=D^{-\frac{1}{2}}{(A^{T}A)e}$ $\times D^{\frac{1}{2}}(A^{T}A)B\quad x_{j}=\frac{\sum_{i}{A_{ij}B_{i}}}{\sum_{i}A_{ij}}$ (7)
The proposed initial solution is given in Equ.7, where $D^{\frac{1}{2}}$ is the diagonal operator containing square roots of the row sums, and $e$ is all-ones vector. The expression on the right-hand side is the element-wise formulation. We now proceed to analyze the optimality of the proposed solution.
$\mathcal{L}(x)$ $=\sum_{i=1}^{R}\left\|\sum_{j=1}^{P}A_{ij}x_{j}-B_{i}\right\|$ (8)
$\mathcal{J}(x)$ $=\sum_{i}\sum_{j}A_{ij}\left\|x_{j}-B_{i}\right\|$ (9)

### v2

@@ Abstract
Feature lifting has emerged as a crucial component in 3D scene understanding, enabling the attachment of rich image feature descriptors (e.g., DINO, CLIP) onto splat-based 3D representations. The core challenge lies in optimally assigning rich general attributes to 3D primitives while addressing the inconsistency issues from multi-view images. We present a unified, kernel- and feature-agnostic formulation of the feature lifting problem as a sparse linear inverse problem, which can be solved efficiently in closed form. Our approach admits a provable upper bound on the global optimal error under convex losses for delivering high quality lifted features. To address inconsistencies and noise in multi-view observations, we introduce two complementary regularization strategies to stabilize the solution and enhance semantic fidelity. Tikhonov Guidance enforces numerical stability through soft diagonal dominance, while Post-Lifting Aggregation filters noisy inputs via feature clustering. Extensive experiments demonstrate that our approach achieves state-of-the-art performance on open-vocabulary 3D segmentation benchmarks, outperforming training-based, grouping-based, and heuristic-forward baselines while producing lifted features in minutes. Our code is available in the GitHub. We provide additional website for more visualization, as well as the video.
1 Introduction
@@
Recent advances in splat representations—such as 3D Gaussian Splatting (3DGS) Kerbl et al. (2023), 2D Gaussian Splatting Huang et al. (2024), and Deformable Beta Splatting Liu et al. (2025)—have enabled real-time, high-fidelity scene rendering by modeling geometry with compact, explicit primitives. These splat-based methods combine differentiable projection, alpha compositing, and efficient visibility-aware rasterization to preserve geometric detail and multi-view consistency, supporting applications ranging from style transfer Liu et al. (2024); Galerne et al. (2025) to scene understanding Guo et al. (2024); Qin et al. (2024); Shi et al. (2024); Jun-Seong et al. (2025); Cheng et al. (2024); Cen et al. (2025); Dou et al. (2024); Gu et al. (2024); Zuo et al. (2025); Peng et al. (2024b). Nevertheless, enriching these primitives with detailed descriptors—such as CLIP, DINO, General ViT, and CNN features—remains challenging due to the inherent difficulty of coherently lifting 2D observations into consistent 3D representations.
@@
To address this challenge, recent efforts have focused on lifting per-pixel semantic descriptors (e.g., CLIP, DINO) onto 3D primitives, enabling semantic tasks such as segmentation and querying. Existing semantic feature lifting methods broadly fall into three categories: training-based optimization Shi et al. (2024); Qin et al. (2024); Zhou et al. (2024); Qiu et al. (2024); Zuo et al. (2025), which pioneered in embedding semantic onto 3D primitives via multi-view training; grouping-based association Wu et al. (2025); Peng et al. (2024b); Gu et al. (2024); Liang et al. (2024), which improved efficiency through feature clustering; and heuristic forward methods Guo et al. (2024); Joseph et al. (2024); Dou et al. (2024); Chacko et al. (2025); Jun-Seong et al. (2025); Cheng et al. (2024), which prioritize speed and directly project semantic features onto 3D primitives.
@@
While existing methods have achieved promising results, a unified theoretical framework for feature lifting remains underexplored, especially concerning the heuristic forward methods that bypass the training process. Building upon prior efforts, we identify several open challenges that limit generalization and theoretical understanding. First, establishing a rigorous mathematical formulation for defining feature lifting could be beneficial to enable formal analysis, generalization, and optimization. Second, in the absence of a formal definition, existing approaches currently lack theoretical guarantees regarding the quality of the lifted features, leaving some uncertainty about their proximity to a globally optimal solution. Third, all aforementioned methods are exclusively focus on ”SAM+CLIP” features and ”3DGS” kernels, which may limit generalization across broader settings. Finally, most prior works do not explicitly account for the inherently noisy nature of collected data, where multi-view inconsistencies and observational noise can introduce ambiguity. Our formulation accounts for these challenges through a theoretically grounded linear inverse framework.
Our key contributions are summarized below:
@@
  • Propose the formulation of feature lifting as a sparse linear inverse problem, where each per-primitive descriptor is recovered via a global consistent solution to the system $AX=B$
@@
As illustrated in Figure 1, our framework inputs precomputed splats, sensor parameters, and dense observations. These are mapped to the feature lifting equation $AX=B$, where the Splats Sensor Matrix ($A$) is derived from the geometry and camera parameters, and the observations form the target vector ($B$). To solve for the feature parameters ($X$), we enhance stability by polarizing alpha values to encourage diagonal dominance (Tikhonov Guidance) and employ Post-Lifting Aggregation to discard inconsistent masks. Our formulation is the first to formalize the feature lifting problem as a general linear inverse problem, and can therefore be applied across diverse splat primitives (e.g., 3DGS, 2DGS, Beta Splats) and dense feature modalities (e.g., CLIP, DINO, ViT, ResNet features). Moreover, it establishes a new state-of-the-art on downstream tasks such as open-vocabulary 3D semantic segmentation, outperforming existing training-based, grouping-based, and heuristic-forward baselines in mIoUs.
Figure 1: Overview of our Feature Lifting Framework. Our pipeline lifts dense 2D feature observations (e.g., MaskCLIP, DINO) onto general 3D splat representations by formulating the task as a sparse linear inverse problem. The Solver incorporates Tikhonov Guidance to ensure numerical stability and Post-Lifting Aggregation to filter noisy inputs. The resulting lifted feature parameters enable high-fidelity downstream tasks, such as open-vocabulary 3D segmentation and localization.
2 Related Work
@@
Joint Training-based Methods—such as LangSplat Qin et al. (2024), LeGaussian Shi et al. (2024), FMGS Zuo et al. (2025), and FeatureSplats Qiu et al. (2024)—jointly optimize scene geometry and high‑dimensional feature embeddings to directly learn splat‑wise semantic descriptors. While effective, end-to-end optimization over large descriptor sets and primitives can be computationally intensive, often requiring significant memory and training time. To alleviate resource demands, these approaches employ descriptor compression (e.g., PCA in LangSplat Qin et al. (2024), quantization in LeGaussian Shi et al. (2024), and deep auto‑encoders in FeatureSplats Qiu et al. (2024)) or reduce the number of primitives (FMGS Zuo et al. (2025)). These strategies help manage resource usage, though they may introduce trade-offs in terms of geometric detail or feature precision.
@@
Grouping-based Methods first extract 2D region or instance masks using models such as SAM Kirillov et al. (2023) and SAM2 Ravi et al. (2024). Following that, each mask is linked to 3D Gaussians via a lightweight training process. Finally, features are aggregated into 3D either by additional optimization or by direct unprojection Wu et al. (2025); Peng et al. (2024b); Gu et al. (2024); Liang et al. (2024); Cen et al. (2025). While these approaches reduce the cost of joint optimization, they still require one to two hours of scene‑specific optimization. Furthermore, these methods depend on SAM’s per-view masks, while powerful, are designed primarily for instance segmentation and may not be directly suited for noisy, dense, pixel-wise feature lifting.
@@
In particular, LAGA Cen et al. (2025) uses per-view SAM masks Kirillov et al. (2023) and trains an affinity model to cluster 3D Gaussian splats, subsequently associating splat clusters across views. However, this approach assumes that feature variance reflects true semantic differences and is therefore view-dependent, introducing a dynamic K-means clustering step that adds complexity and may increase memory and computational requirements. In contrast, our analysis suggests that such discrepancies often stem from mask inaccuracies—such as one mask isolating only the target while another over-segments and includes adjacent objects—rather than from genuine viewpoint changes. Our convex-regularized inverse problem solver suppresses this noise directly—without requiring additional training, multi-level clustering algorithms, or view-dependent features assigned to one primitive.
@@
Heuristic Forward Methods-e.g., Argmax Lifting Chacko et al. (2025), Occam’s LGS Cheng et al. (2024), Semantic Gaussian Guo et al. (2024), gradient-guided splitting Joseph et al. (2024), CosegGaussians Dou et al. (2024), and DrSplats Jun-Seong et al. (2025)—are analytic and training-free, offer high efficiency and simplicity. However, they often lack a formal mathematical foundation and sensitive to noisy inputs. Interestingly, three recent heuristic pipelines—CosegGaussians Dou et al. (2024), Occam’s LGS Cheng et al. (2024), and DrSplats Jun-Seong et al. (2025)—independently proposed the identical row-sum weighting rule (row-sum preconditioner), underscoring its broad practical effectiveness. However, while CosegGaussians applied the row‐sum weighting scheme $x_{i}=\sum_{j}w_{i,j}b_{j}/\sum_{j}w_{i,j}$ , it did so without an accompanying theoretical interpretation. Occam’s LGS later offered a maximum‐likelihood interpretation of the same weighted‐average rule but provided no error bound. Dr. Splat simplifies the preconditioner further by summing only the top-$k$ contributions per primitive, trading off some theoretical rigor for computational efficiency.
@@
3 Feature Splat Solver
@@
3.1 Problem Definition
3.1.1 Splats Rendering
To justify formulating the Feature Lifting problem as a linear inverse problem, we first review the core splats rendering process. Most splats primitives, including planar 2D Gaussian Splats Huang et al. (2024), volumetric 3D Gaussian Splats Yu et al. (2024); Kerbl et al. (2023), and Deformable Beta Splats Liu et al. (2025), employ a fixed, depth-sorted alpha-blending pipeline. Concretely, for each viewing ray $r$, there is a sequence of primitives ordered from front to back, with their respective contribution to the final color denoted as $\omega_{p}$. The background color is $C_{b}$. The rendered ray color $C_{r}$ is given below:
@@ 3.1.1 Splats Rendering
The rendered color $C_{r}$ is usually very close (i.e. PSNR is more than 24 according to Yu et al. (2024); Kerbl et al. (2023); Liu et al. (2025)) to the given observation $\hat{C}_{r}$ in the training data set.
The first property arises from the tile-based rendering design of Gaussian SplatsKerbl et al. (2023). Since the tile size is typically set to $16\times 16$, most splats are excluded from rendering on any given tile.
The second property is the row stochastic property. It is less than one because the standard alpha‐blending heuristic of terminating ray integration once the composite opacity approaches one. It is close to one based on the assumption that scene geometry should block any further contribution from background color during training. In practice, we randomize the background color during training, forcing splats to fully occlude the randomly colored background from all viewing directions. Further justification of this property can be found in the appendix.G.
The third observation is noted in a recent splats training method Yu et al. (2024); Huang et al. (2024); Kerbl et al. (2023); Liu et al. (2025). Those properties will be used to justify our formulation.
3.1.2 Feature Lifting Equation
@@
3.1.3 Inverse Problem
First, we ask whether Equ.3 admits a solution, and, whether the given solution is unique. In typical settings—where $A\in R^{R\times P}$ with $R\gg P$— the system is overdetermined and generally inconsistent, meaning no exact $X$ satisfies equation 3. Instead, we consider the least squares formulation (or more generally, a convex loss minimization problem):
@@ 3.1.3 Inverse Problem
where the minimizer exists for any $B$ and lies in the column space of $A$. In our context, Property 3 guarantees that real‐world observations $B$ , such as RGB color, approximately lie within $\mathrm{range}(A)$. As a result, the least-squares residual is small, and a near-exact lift exists. Intuitively, any observation derived from the same RGB signal should have a near-exact lift. From a semantic viewpoint, we also require a one‐to‐one mapping. Each geometric primitive should admit a single descriptor, mirroring an ideal embedding (e.g. a perfect 3D CLIP) that assigns one feature per object. Thus, uniqueness in equation 5 is not just for mathematical convenience, but for the purpose of aligning with downstream task requirements. Any deviation from uniqueness signals is either due to noisy inputs or suboptimal solvers—precisely the singular, and misleading-mask problem we address through our convex regularization and preconditioning strategies.
@@
3.2 Solver
In this session, we will introduce our solver under the well-posedness condition and give an optimal upper bound of our solver. In our setting, the least‐squares problem in Eq. equation 5 is convex, so stochastic methods such as SGD provably converge to the global minimizer Esser et al. (2010); Saad (2003). However, cold‐start training can be prohibitively slow. Prior works (e.g. Chacko et al. (2025); Dou et al. (2024); Jun-Seong et al. (2025); Cheng et al. (2024); Joseph et al. (2024)) have explored efficient one-shot approximations that show practical value but do not offer formal guarantees. To address this, we introduce the row‐sum preconditioner as a, closed‐form solver. We then develop an analytical framework that (i) quantifies its $(1+\beta)$-approximation error under both L2 and general convex losses, and (ii) interprets the behavior of existing heuristics as special cases.
$D^{\frac{1}{2}}(A^{T}A)=\sqrt{D(A^{T}A)}\quad x=D^{-\frac{1}{2}}{(A^{T}A)e}\times D^{\frac{1}{2}}(A^{T}A)B\quad x_{j}=\frac{\sum_{i}{A_{ij}B_{i}}}{\sum_{i}A_{ij}}$ (6)
The proposed initial solution is given in Equ.6, where $D^{\frac{1}{2}}$ is the diagonal operator containing square roots of the row sums, and $e$ is all-ones vector. The expression on the right-hand side is the element-wise formulation. We now proceed to analyze the optimality of the proposed solution.
$\mathcal{L}(x)=\sum_{i=1}^{R}\left\|\sum_{j=1}^{P}A_{ij}x_{j}-B_{i}\right\|$ $\quad\mathcal{J}(x)=\sum_{i}\sum_{j}A_{ij}\left\|x_{j}-B_{i}\right\|$ (7)